The perimeter of a parking lot is 310 meters. The length is 10 more than twice the width. Find the length and width.
Answer by fractalier(6550)

We have perimeter P = 2L + 2W = 310 and
L = 2W + 10
Substitute this into the first equation and get
2(2W + 10) + 2W = 310
4W + 20 + 2W = 310
6W = 290
W = 48 1/3 m
L = 2(48 1/3) + 10 = 106 2/3 m